OpenStudy (griffindore):

Sue buys a certain brand of cereal that costs $12 per box. Sue changes to a super-saving brand of the same size. The equation shows the price, y, as a function of the number of boxes, x, for the new brand. y = 8x Part A: How many more dollars is the price of a box of Sue's original brand of cereal than the price of a box of the super-saving brand? Show your work. Part B: How much money does Sue save each month with the change in cereal brand if she buys 4 cereal boxes each month? Show your work.

OpenStudy (griffindore):

Do i multiply or subtract?

OpenStudy (welshfella):

what is the price of 1 box of the super-saving cereal?

OpenStudy (griffindore):

8

OpenStudy (welshfella):

right 8 dollars So how many more dollars was one box of the original brand?

OpenStudy (griffindore):

4

OpenStudy (welshfella):

correct 12 - 8 = 4

OpenStudy (griffindore):

oh ok

OpenStudy (welshfella):

So any ideas about Part B?

OpenStudy (griffindore):

umm not realy

OpenStudy (welshfella):

How much does she spend each month buying new cereal and how much buying the original?

OpenStudy (griffindore):

8

OpenStudy (welshfella):

8? ONE box costs $8

OpenStudy (welshfella):

- how many does you buy each month?

OpenStudy (welshfella):

- it says in the question...

OpenStudy (griffindore):

sorry

OpenStudy (griffindore):

oh 32

OpenStudy (welshfella):

4 boxes in a month so cost = 8*4 = $32 Now work out the cost for the original cereal ( 4 at $12) and subtract $32 dollars from this and you have the answer

OpenStudy (griffindore):

ok

OpenStudy (welshfella):

OK? can you do that?

OpenStudy (griffindore):

i did it

OpenStudy (griffindore):

its 98

OpenStudy (welshfella):

???? No what is the cost of 4 boxes of the original cereal?

OpenStudy (griffindore):

oh 128

OpenStudy (griffindore):

but the difference is 98

OpenStudy (welshfella):

cost = number of boxes * price of each box = 4 * 12 = ?

OpenStudy (griffindore):

128

OpenStudy (welshfella):

no 4 * 12 = 4*10 + 4 * 2

OpenStudy (griffindore):

oh crap its 48

OpenStudy (welshfella):

right!! so savings = 48 - 32

OpenStudy (griffindore):

idk why i thought it was 128

OpenStudy (griffindore):

which is 16

OpenStudy (welshfella):

nor me!! lol!

OpenStudy (griffindore):

lol

OpenStudy (welshfella):

16 is the answer Dont forget to show all your work.

OpenStudy (griffindore):

i did

OpenStudy (welshfella):

good

OpenStudy (griffindore):

thx i have another one, but im just gonna post it instead of putting it here